OpenStudy (prepfortears):

Make a perfect square trinomial: 25^2+__x+16

OpenStudy (danjs):

square(one length sides) (x + a)*(x+a) = (x + a)^2

OpenStudy (danjs):

root 25 = 5 root 16 = 4 you want to add the value that corresponds to the expansion of (a*x + b)^2 = (a*x + b) * (a*x + b)

OpenStudy (danjs):

a^2*x^2 + 2*(a*b)*x + b^2

OpenStudy (danjs):

You have 25*x^2+ __ x + 16 compare that to the expansion above a^2*x^2 + 2*(a*b)*x + b^2 so, a=5 , b = 4, the middle value is 2*5*4 or 40